# VILA-U: a Unified Foundation Model Integrating Visual Understanding and Generation

### v2

@@ 4.3 Quantitative Evaluation
LLaVA-1.5 llava Vicuna-1.5-7B Continuous 336 78.5 62.0 58.2 85.9 1510.7 58.6 30.5
VILA lin2023vila LLaMA-2-7B Continuous 336 79.9 62.3 64.4 85.5 1533.0 61.1 34.9
Unified-IO 2 Unified-io2 6.8B from scratch Continuous 384 79.4 87.7 61.8
InstructBLIP instructblip Vicuna-7B Continuous 224 49.2 50.1 53.4 26.2
IDEFICS-9B idefics LLaMA-7B Continuous 224 50.9 38.4 25.9
Emu Emu LLaMA-13B Continuous 224 52.0
LaVIT jin2023unified LLaMA-7B Continuous 224 66.0 46.8
DreamLLM dreamllm Vicuna-7B Continuous 224 72.9 41.8 36.6
Video-LaVIT video-lavit LLaMA-2-7B Continuous 224 80.2 63.6 1581.5 64.4 35.0
CM3Leon-7B CM3Leon 7B from scratch Discrete 256 47.6
LWM lwm LLaMA-2-7B Discrete 256 55.8 44.8 18.8 75.2 9.6
Show-o xie2024show Phi-1.5-1.3B Discrete 256 59.3 48.7 73.8 948.4
@@
Unified-IO 2 Unified-io2 6.8B from scratch Continuous 384 52.1 42.5
Emu Emu LLaMA-13B Continuous 224 18.8 8.3
VideoChat videochat Vicuna-7B Continuous 224 56.3 45 34.4
Video-LLaMA video-llama LLaMA-2-7B Continuous 224 51.6 29.6
Video-ChatGPT video-chatgpt LLaMA-2-7B Continuous 224 64.9 49.3 51.4 35.2
Video-LLava video-llava Vicuna-7B Continuous 224 70.7 59.2 70.0 45.3
Video-LaVIT video-lavit LLaMA-2-7B Continuous 224 73.5 59.5 50.2
LWM lwm LLaMA-2-7B Discrete 256 55.9 44.1 40.9
@@
SD-XL sdxl Diffusion 2000M 9.55
PixArt chen2023pixart Diffusion 25M 6.14
Playground v2.5 playgroundv2.5 Diffusion 4.48
Show-o xie2024show Discrete Diff. 36M 15.18
LWM lwm Autoregressive 17.77
@@
Table 5: Comparison with other visual generation methods on GenAI-Bench GenAI-Bench . The results show that our method outperforms previous autoregressive visual generation methods. For advanced prompts that require better text following ability to generate, our method can have a relatively small performance gap with diffusion-based methods, even with much less training data and time.
@@
SD v2.1 ldm Diffusion 2000M 0.80 0.79 0.76 0.77 0.80 0.78
SD-XL sdxl Diffusion 2000M 0.84 0.84 0.82 0.83 0.89 0.83
Midjourney v6 midjourney Diffusion – 0.88 0.87 0.87 0.87 0.91 0.87
DALL-E 3 dalle-3 Diffusion – 0.91 0.90 0.92 0.89 0.91 0.90
Show-o xie2024show Discrete Diff. 36M 0.72 0.72 0.70 0.70 0.75 0.70
LWM lwm Autoregressive – 0.63 0.62 0.65 0.63 0.70 0.63
@@
SD v2.1 ldm Diffusion 2000M 0.68 0.70 0.68 0.54 0.64 0.62
SD-XL sdxl Diffusion 2000M 0.71 0.73 0.69 0.50 0.66 0.63
Midjourney v6 midjourney Diffusion – 0.78 0.78 0.79 0.50 0.76 0.69
DALL-E 3 dalle-3 Diffusion – 0.82 0.78 0.82 0.48 0.80 0.70
Show-o xie2024show Discrete Diff. 36M 0.70 0.62 0.71 0.51 0.65 0.60
LWM lwm Autoregressive – 0.59 0.58 0.54 0.49 0.52 0.53
@@
Visual Generation Tasks. As shown in Table 4, VILA-U can achieve a better FID than other autoregressive methods and have comparable performance with some diffusion based methods. This result shows the feasibility of our method for visual generation. Table 4.3 summarizes the quantitative results of our method and other visual generation methods on GenAI-Bench. Although Our method is inferior to diffusion-based visual generation methods that have been trained on billions-level image-text pairs, our method has comparable performance with SD v2.1 rombach2022high and SD-XL sdxl on advanced prompts even trained with magnitude-level less data. This further shows that VILA-U can learn the correlation among visual and textual modalities effectively and efficiently with our unified training framework.
@@ 4.4 Qualitative Evaluation
Visual Understanding. To validate the effectiveness of VILA-U in comprehensive visual understanding tasks, we apply it in several understanding and reasoning tasks, as some examples shown in Figure 3 and Figure 5. From the results, we can see the versatility of VILA-U in various tasks including visual captioning and visual question answering. Besides, our model has inherited some important capabilities from VILA lin2023vila including multi-image understanding, in-context learning, as shown in Figure 5 and Figure 6.
@@
Visual Generation. We present some examples of the visual generation results with VILA-U in Figure 7. We can see that our model can be employed in both image generation and video generation, even trained with a relatively small data corpus. In the given examples, our method can generate nice-looking images and continuous videos adhering to users’ input.
@@ 5 Ablation Study
In this section, we delve into some key design components of our framework and design ablation experiments to show how they will influence the performance.
@@ 5.2 Impact of Contrastive Loss to Visual Generation
We conduct two experiments to demonstrate the influence of contrastive loss to generation performance. For efficiency, we conduct only text-to-image pretraining and utilize Sheared-LLaMA-1.3B xia2023sheared instead of LLaMA-2-7B as the LLM. In the first experiment, we use the RQ-VAE as the vision tower, which has an rFID of 1.30. In the second experiment, we employ our unified vision tower. Results are shown in Table 5.2. Our Unified Vision Tower yielded slightly worse FID results than the RQ-VAE on MJHQ-30K, possibly due to its inferior rFID resulting from the introduction of contrastive loss.
@@
RQ-VAE lee2022autoregressive Sheared-LLaMA-1.3B 256 $\times$ 256 1.30 12.0
@@ 6 Conclusion
We present VILA-U, a novel and unified visual language model that integrates video, image and language understanding and generation tasks into one autoregressive next-token prediction framework. Our method is not only more concise than most VLMs that leverage additional components like diffusion models for unifying visual generation and understanding, but also demonstrates that autoregressive methods can achieve comparable performance to state-of-the-art VLMs. Our success is due to both a unified foundation vision tower that aligns discrete visual features with texts during pre-training and a high-quality dataset suitable for visual understanding and generation training. We believe VILA-U can serve as a general-purpose framework for diverse visual language tasks.
Limitations. There is still a performance gap in the visual understanding ability between VILA-U and state-of-the-art VLMs leveraging continuous visual feature. Besides, the visual generation quality is relatively low compared to state-of-the-art diffusion models. In future work, we will be committed to overcoming these limitations to build an advanced VLM that can achieve state-of-the-art performance in all kinds of visual language tasks.

### v3

@@ 4.3 Quantitative Evaluation
LLaVA-1.5 Vicuna-1.5-7B Continuous 336 78.5 62.0 58.2 85.9 1510.7 58.6 30.5
VILA LLaMA-2-7B Continuous 336 79.9 62.3 64.4 85.5 1533.0 61.1 34.9
Unified-IO 2 6.8B from scratch Continuous 384 79.4 87.7 61.8
InstructBLIP Vicuna-7B Continuous 224 49.2 50.1 53.4 26.2
IDEFICS-9B LLaMA-7B Continuous 224 50.9 38.4 25.9
Emu LLaMA-13B Continuous 224 52.0
LaVIT LLaMA-7B Continuous 224 66.0 46.8
DreamLLM Vicuna-7B Continuous 224 72.9 41.8 36.6
Video-LaVIT LLaMA-2-7B Continuous 224 80.2 63.6 1581.5 64.4 35.0
Emu2-Chat Emu2-37B Continuous 448 84.9 65.1 66.6
MM-Interleaved Vicuna-13B Continuous 224 80.2 60.5 61.0
DEEM Vicuna-7B Continuous 448 68.2 55.7 37.4
CM3Leon-7B 7B from scratch Discrete 256 47.6
LWM LLaMA-2-7B Discrete 256 55.8 44.8 18.8 75.2 9.6
Show-o Phi-1.5-1.3B Discrete 256 59.3 48.7 73.8 948.4
SEED-LLaMA Vicuna-7B Discrete 224 66.2 51.5
@@
Unified-IO 2 6.8B from scratch Continuous 384 52.1 42.5
Emu LLaMA-13B Continuous 224 18.8 8.3
VideoChat Vicuna-7B Continuous 224 56.3 45 34.4
Video-LLaMA LLaMA-2-7B Continuous 224 51.6 29.6
Video-ChatGPT LLaMA-2-7B Continuous 224 64.9 49.3 51.4 35.2
Video-LLava Vicuna-7B Continuous 224 70.7 59.2 70.0 45.3
Video-LaVIT LLaMA-2-7B Continuous 224 73.5 59.5 50.2
Emu2-Chat Emu2-37B Continuous 448 49.0 31.4
LWM LLaMA-2-7B Discrete 256 55.9 44.1 40.9
SEED-LLaMA Vicuna-7B Discrete 224 40.9 30.8
@@
SD v2.1 Diffusion 26.96
SD-XL Diffusion 2000M 9.55
PixArt Diffusion 25M 6.14
Playground v2.5 Diffusion 4.48
LWM Autoregressive 17.77
Show-o Autoregressive 36M 15.18
@@
Table 5: Comparison with other visual generation methods on GenAI-Bench (Lin et al., 2024). The results show that our method outperforms previous autoregressive visual generation methods. For advanced prompts that require better text following ability to generate, our method can have a relatively small performance gap with diffusion-based methods, even with much less training data.
@@
SD v2.1 Diffusion 2000M 0.80 0.79 0.76 0.77 0.80 0.78
SD-XL Diffusion 2000M 0.84 0.84 0.82 0.83 0.89 0.83
Midjourney v6 Diffusion – 0.88 0.87 0.87 0.87 0.91 0.87
DALL-E 3 Diffusion – 0.91 0.90 0.92 0.89 0.91 0.90
LWM Autoregressive – 0.63 0.62 0.65 0.63 0.70 0.63
Show-o Autoregressive 36M 0.72 0.72 0.70 0.70 0.75 0.70
@@
SD v2.1 Diffusion 2000M 0.68 0.70 0.68 0.54 0.64 0.62
SD-XL Diffusion 2000M 0.71 0.73 0.69 0.50 0.66 0.63
Midjourney v6 Diffusion – 0.78 0.78 0.79 0.50 0.76 0.69
DALL-E 3 Diffusion – 0.82 0.78 0.82 0.48 0.80 0.70
LWM Autoregressive – 0.59 0.58 0.54 0.49 0.52 0.53
Show-o Autoregressive 36M 0.70 0.62 0.71 0.51 0.65 0.60
@@
Visual Generation Tasks. As shown in Table 4, VILA-U can achieve a better FID than other autoregressive methods and have comparable performance with some diffusion based methods. This result shows the feasibility of our method for visual generation. Table 4.3 summarizes the quantitative results of our method and other visual generation methods on GenAI-Bench. Although Our method is inferior to diffusion-based visual generation methods that have been trained on billions-level image-text pairs, our method has comparable performance with SD v2.1 (Rombach et al., 2022b) and SD-XL (Podell et al., 2023) on advanced prompts even trained with magnitude-level less data. This further shows that VILA-U can learn the correlation among visual and textual modalities effectively with our unified training framework. For video generation, we evaluate our method on VBench (Huang et al., 2024) and compare it against Open-Sora (Zheng et al., ), CogVideo (Hong et al., 2022), and CogVideoX (Yang et al., 2024). The results, presented in Table 6, demonstrate that our method achieves performance that is better than CogVideo and comparable to Open-Sora, highlighting the effectiveness of our approach.
[…]
Open-Sora 75.91 78.82 64.28
CogVideo 67.01 72.06 46.83
CogVideoX 81.61 82.75 77.04
Ours (256) 74.01 76.26 65.04
Table 6: Comparison with other visual generation methods on VBench (Huang et al., 2024).
@@ 4.4 Qualitative Evaluation
Visual Understanding. To validate the effectiveness of VILA-U in comprehensive visual understanding tasks, we apply it in several understanding and reasoning tasks, as some examples shown in Figure 3 and Figure 5. From the results, we can see the versatility of VILA-U in various tasks including visual captioning and visual question answering. Besides, our model has inherited some important capabilities from VILA (Lin et al., 2023) including multi-image understanding, in-context learning, as shown in Figure 5 and Figure 6. More visualizations can be found in the Appendix B.2 and B.3.
@@
Visual Generation. We present some examples of the visual generation results in Figure 7. Our model can be employed in both image generation and video generation, even trained with a relatively small data corpus. In the given examples, our method can generate nice-looking images and continuous videos adhering to the user’s input. More visualizations can be found in the appendix B.4.
@@ 5.2 Impact of Contrastive Loss to Visual Generation
We conduct two experiments to demonstrate the influence of contrastive loss to generation performance. For efficiency, we conduct only text-to-image pretraining and utilize Sheared-LLaMA-1.3B (Xia et al., 2023) instead of LLaMA-2-7B as the LLM. In the first experiment, we use the RQ-VAE as the vision tower, which has an rFID of 1.30. In the second experiment, we employ our unified vision tower. Results are shown in Table 5.2. Our Unified Vision Tower yielded slightly worse FID results than the RQ-VAE on MJHQ-30K, possibly due to its inferior rFID resulting from the contrastive loss.
@@
RQ-VAE (Lee et al., 2022) Sheared-LLaMA-1.3B 256 $\times$ 256 1.30 12.0
@@ 6 Conclusion and Limitation
We present VILA-U, a novel and unified visual language model that integrates video, image and language understanding and generation tasks into one autoregressive next-token prediction framework. Our method is not only more concise than most VLMs that leverage additional components like diffusion models for unifying visual generation and understanding, but also demonstrates that autoregressive methods can achieve comparable performance to state-of-the-art VLMs. We believe VILA-U can serve as a general-purpose framework for diverse visual language tasks.
As demonstrated in Section 5.2, the introduction of contrastive loss impacts the reconstruction ability of the vision tower. Balancing these two capabilities within the unified vision tower presents an interesting and complex challenge that requires further exploration. Additionally, we currently do not observe significant synergy or mutual enhancement between understanding and generation tasks. In the future, we aim to investigate and explore more effective methods to enable these tasks to complement and reinforce each other, thereby fully realizing the untapped potential of a unified visual language model.
Appendix
Appendix A Difference with related works
Prior to VILA-U, unified visual language models were dominated by two mainstream approaches:
(1) Represented by LWM, CM3Leon and Show-o which utilizes a VQGAN-based tokenizer to convert visual inputs into discrete tokens. However, as these tokenizers are trained solely with a reconstruction objective, the resulting tokens lack rich semantic information. This limitation leads to poor performance on multimodal understanding tasks. But it can easily support autoregressive visual generation and the generated visual tokens can be seamlessly decoded into visual outputs using the lightweight decoder of VQGAN.
(2) Represented by AnyGPT SEED-LLaMa and LaViT, which utilizes a codebook to quantize features produced by a pre-trained ViT model like CLIP. Since CLIP features encode rich semantic information, these approaches generally achieve significantly better performance on understanding tasks compared to VQGAN-based tokenizers. However, these tokenizers lack decoding capability, requiring an external visual generation model, such as a diffusion model, to use the generated visual tokens as conditions for producing visual outputs.
Compared to these two mainstream approaches, VILA-U introduces a solution that addresses the limitations of both. We design a unified vision tower that extracts features with rich semantic information, similar to CLIP, while also supporting image reconstruction capabilities akin to VQGAN. This is achieved by incorporating both reconstruction loss and contrastive loss into the autoencoder training process, along with utilizing residual quantization to enhance the representation capability of the visual features. Building on this foundation, we develop a single end-to-end autoregressive framework that eliminates the need for external visual generation models required by approach 2 and significantly outperforms the understanding results of methods in approach 1.
Appendix B Qualitative Results
B.1 Reconstruction
Figure 8: Visualization of the reconstruction results from text-aligned discrete visual tokens.
We present qualitative reconstruction results in Figure 8 for our 256 / 384 resolution vision tower. These vision towers effectively reconstruct images in detail using text-aligned discrete visual tokens.
B.2 Visual Understanding
Figure 9: Image understanding results. Examples are taken from the test split of VQAv2 dataset.
Figure 10: Video understanding results. Examples are taken from the test split of TGIF dataset.
We provide more image understanding and video understanding examples in Figure 9 and Figure 10. VILA-U successfully answers the questions accurately.
B.3 In-context Learning Examples
Figure 11: In-context learning examples. We try all in-context learning examples in Lin et al. (2023). The results demonstrate that VILA-U has inherited good in-context learning capabilties.
We provide more qualitative results to demonstrate in-context learning capabilities of VILA-U in Figure 11. VILA-U exhibits good in-context learning capabilties.
B.4 Visual Generation
Figure 12: Image generation results. VILA-U can generate high-quality images given text input.
Figure 13: Video generation results. VILA-U can generate high-quality videos given text input.
We provide more image generation and video generation examples in Figure 12 and Figure 13. VILA-U can generate high-quality images and videos given text input.
Appendix C Failed Training Recipes.
We experiment with numerous training recipes and find none to be as effective as our final approach. We list four alternative recipes and discuss their shortcomings compared to our final recipe: 1) Load pre-trained CLIP weights into the text encoder only; 2) Load pre-trained RQ-VAE weights for the vision encoder and decoder while training other parts from scratch; 3) Freeze the vision encoder; 4) Make the text encoder trainable.